OpenStudy (fibonaccichick666):

Find the limit without using L'Hopital's rule: \[lim_{h\rightarrow 0} \frac{21.7^h-1}{h}\]

OpenStudy (fibonaccichick666):

Took the ln of both sides. to separate, then took the limit, but I keep getting zero when I need 3.07731

OpenStudy (fibonaccichick666):

@e.mccormick

OpenStudy (fibonaccichick666):

if you could offer your valuable insight pretty please

OpenStudy (e.mccormick):

OK, I can see where the 0 is coming from.... https://www.desmos.com/calculator/f95qx89kca

OpenStudy (fibonaccichick666):

right, but i need 3.blahblah

OpenStudy (e.mccormick):

Did you try separating the fraction into two fractions?

OpenStudy (fibonaccichick666):

didn't help, I get something I can't apply a log to either

OpenStudy (fibonaccichick666):

and evaluating the limit i get the first fraction minus infinity, which is not helpful

OpenStudy (e.mccormick):

Yah, I was just noticing that... which is worse than getting 0!

OpenStudy (fibonaccichick666):

Idk, where I went wrong lol indeed worse

OpenStudy (fibonaccichick666):

i have this, \[ \lim_{h\rightarrow 0} ln(21.7^h-1)-ln(h)=ln(L)\]

OpenStudy (fibonaccichick666):

after that take the lim, still get 0

OpenStudy (fibonaccichick666):

you actually get 1 but

OpenStudy (fibonaccichick666):

still wrong answer

OpenStudy (anonymous):

you may want to mult top and bottom by \(27.1^h+1\)

OpenStudy (fibonaccichick666):

ok let me give that a shot,

OpenStudy (fibonaccichick666):

nah, not much help I still can't get rid of the minus 1

OpenStudy (anonymous):

what about the squeeze theorem?

OpenStudy (fibonaccichick666):

nope, I don't know the answer from the get go

OpenStudy (anonymous):

rewrite the limit.\[\lim_{h \rightarrow 0}\frac{ 27.1^h-1 }{ h }=\lim_{h \rightarrow 0}\frac{ e^{h\ln(27.1)}-1 }{ h }=\ln(27.1)\cdot \lim_{h \rightarrow 0}\frac{ e^{h\ln(27.1)}-1 }{ \ln(27.1)\cdot h }\]now let \(k=\ln(27.1)\cdot h\) then as \(h\rightarrow 0, k\rightarrow 0\) and you get \[\ln(27.1)\cdot \lim_{h \rightarrow 0}\frac{ e^{h\ln(27.1)}-1 }{ \ln(27.1)\cdot h }=\ln(27.1)\cdot \lim_{k \rightarrow 0}\frac{ e^{k}-1 }{ k }=\ln(27.1)\]

OpenStudy (anonymous):

hey @FibonacciChick666 , have a look

OpenStudy (fibonaccichick666):

hmm last step, that's 0/0 l'hops gives you e^k-1=0 so stilll get 0 right?

OpenStudy (fibonaccichick666):

i like the e^ln idea though

OpenStudy (anonymous):

no, \[\lim_{h \rightarrow 0}\frac{ e^h-1 }{ h }=1\]by definition

OpenStudy (fibonaccichick666):

really? I must be spacing sorry, how?

OpenStudy (fibonaccichick666):

oh nvm i'm an idiot

OpenStudy (anonymous):

you good?

OpenStudy (fibonaccichick666):

yea ok, that works thank you

OpenStudy (fibonaccichick666):

Idk why I couldn't think of anything

OpenStudy (anonymous):

sorry, i forgot that anytime you have another base you always convert to base e (same goes for logs... always convert to ln).

OpenStudy (fibonaccichick666):

yea, I got that, I just couldn't think of how to do this without l'hops

OpenStudy (fibonaccichick666):

they haven't learned it yet

OpenStudy (anonymous):

you can always check Paul's Online Math Notes. He's even got some nice "cheat sheets" there too!

OpenStudy (fibonaccichick666):

yea, I know it. I'm just helping my bf, and for the life of me, couldn't figure this one out... calc was like 4 years ago O.O

OpenStudy (anonymous):

have fun!

OpenStudy (fibonaccichick666):

indeed. thanks again!